If block B weighs 200 lb and block C weighs 100 lb
Answer #1

let C weight be W
tension in C string=W
tension in B string=275
for equilibrium at A
balancing horizontal forces
Wcos30=275cosθ

balacing vertical forces

Wsin30+275sinθ=300

275cosθtan30+275sinθ=300

solving for θ we get

cosθ=0.188

θ=79.13 degrees

W=275*0.188/cos30

=59.7 lb
Answer #2

uploaded imageuploaded imageuploaded imageuploaded image